OpenStudy (anonymous):

Write the equation of an ellipse with foci at (-4, 0) and (-4, -6) and major axis of 10

OpenStudy (anonymous):

is it (y+3)^2/25+(x+4)^2/16=1? @jim_thompson5910

OpenStudy (anonymous):

And this one: Convert the following equation from rectangular to polar form: x = 2

jimthompson5910 (jim_thompson5910):

(y+3)^2/25+(x+4)^2/16=1 is correct

jimthompson5910 (jim_thompson5910):

x = 2 r*cos(t) = 2 r = 2/cos(t) r = 2*sec(t)